Solve each of the following systems by substitution. I am having major problems with these kind of problems...HELP, ZACK
5x - 2y= -5
1 y - 5x = 3
8x - 4y = 16
y = 2x - 4
This last problem the 4y and the y in the next prblem line up buy it isnt turning out like that before i submit it

Found 2 solutions by brianuh, MathTherapy:
Answer by brianuh(2)

5x - 2y= -5

Divide everything by 2y to get Y by itself.
so 5x divided by -2
y = -5/2x 5/2
the -5 becomes postive because negative 5 divided by negative two is a positive, so positive 5 over 2.
the answer is
y = -5/2x + 5/2




8x - 4y = 16
Divide by -4y on all sides to get y by itself. That is the ultimate goal, y to be by itself.
8 divided by -4 is -2x and 16 divided by -4 is -4.
so y = -2x -4

Answer by MathTherapy(10858)

Solve each of the following systems by substitution.  I am  having major problems with these kind of problems...HELP,  ZACK

5x - 2y= -5
1 y - 5x = 3

8x - 4y = 16
      y = 2x - 4

I guess , by you saying that ".....This last problem the 4y and the y in the next prblem line up buy it isnt turning out like
that before i submit it
***********************
The following response by the other person, is mere RUBBISH!! It's so senseless that this author thinks that a person who knows nothing
about math, will agree!  

"5x - 2y= -5

Divide everything by 2y to get Y by itself.
so 5x divided by -2
y = -5/2x 5/2
the -5 becomes postive because negative 5 divided by negative two is a positive, so positive 5 over 2.
the answer is
y = -5/2x + 5/2

8x - 4y = 16
Divide by -4y on all sides to get y by itself. That is the ultimate goal, y to be by itself.
8 divided by -4 is -2x and 16 divided by -4 is -4.
so y = -2x -4"
==============
For the 1st system, see: Equations/94042 for this author's response.

2nd system:  8x - 4y = 16___2x - y = 4___2x - 4 = y ---- eq (i)
                                 y = 2x - 4 ---- eq (ii)

As seen above, both equations are EXACTLY the SAME, so this constitutes a CONSISTENT/DEPENDENT SYSTEM, with an INFINITE number
of solutions, as both equations represent the same line. This author believes this is why you may've had problems solving it.